OpenStudy (anonymous):

During a soccer game, Erin scored 1 out of 2 goals attempted. Lauren scored 6 out of 8 goals attempted. Write the fraction that represents each and compare. Who was the most successful?

OpenStudy (anonymous):

*Erin; 3/4* Lauren- 2/5

OpenStudy (anonymous):

erin was more successful:)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Actually I think Lauren was more successful. Because Erin scored only half (50%) of his attempts, but Lauren scored 75% of her attempts. So Lauren was more successful.

OpenStudy (anonymous):

I'm so confused

OpenStudy (anonymous):

im pretty sure erin was still more successful because erin got 1 out of 2 and laurn got 6 out of 8...i think your percentages might be mixed up?

OpenStudy (anonymous):

erin only missed one and laurn missed 2

OpenStudy (anonymous):

Well, you need to look at this problem from this perspective. Say Erin scored 1 out of 2 goals that he attempted. The fraction that represents this is,\[\frac{1}{2} = 0.5\]Now, this means he was 50% correct in his attempts. Now, Lauren shot 8 times, but scored 6 times. This means we write this fraction:\[\frac{6}{8} = \frac{3}{4}=0.75\]This means she was 75% correct in her attempts. I think that it doesn't matter how much Erin or Lauren missed, what matters is how much they scored with respect to their attempts. We know that,\[0.75 > 0.5\]

OpenStudy (anonymous):

oh..ok;)